A roller-coaster car with a mass of 500kg moves at the bottom of a circular dip of radius 40.0m with a speed of 10.0m/s. Find the normal force of the track on the car at the bottom of the dip. 



1
Expert's answer
2021-05-27T12:29:11-0400

Mass(m)=500kg

Radius (r)=40 m

Speed(v)=10.00m/sec

Force=?

"a_c=\frac{v^2}{r}"

Put value

"a_c=\frac{10^2}{40}=\frac{100}{40}=2.5m/sec^2"

Force

F=mac

Put value

"F=500\times2.5N"

"F=1250N"